OpenStudy (anonymous):

"Solve" A piece of rope 5m long is cut into two pieces so that one piece is three fifths as long as the other. Find the length of each piece.

OpenStudy (anonymous):

so you have two pieces, pretend one is X meters long. How long is the other piece in terms of X?

OpenStudy (anonymous):

3/5? idk lol im really bad at algebra 2

OpenStudy (anonymous):

No you're not, that's correct!

OpenStudy (anonymous):

OH YAY!! :D

OpenStudy (anonymous):

Now, you know one piece is X meters, and the other is 3/5 * X meters. and the total is 5 meters. Can you write an equation to solve?

OpenStudy (anonymous):

um x+3/5=5?

OpenStudy (anonymous):

Close, you don't want to lose the other X. it's \[x + (3/5)*x = 5 meters\]

OpenStudy (anonymous):

OHH forgot the other x sorry

OpenStudy (anonymous):

No need to apologize, just recognize :D Now, do you know how to simplify the left hand side?

OpenStudy (anonymous):

haha I've been bad at algebra 1 and 2 lol.. do i -x or is it 2x? im not sure

OpenStudy (anonymous):

you need to combine the two x terms. You have 1 X plus 3/5 X, how many X is that total?

OpenStudy (anonymous):

dang fraction lol that is what i dont get.. i need help on that one

OpenStudy (anonymous):

If it helps, you can pretend the Xes aren't there. It's 1 (aka 5/5) plus 3/5 for a total of 8/5. So the equation is 8/5*x = 5

OpenStudy (anonymous):

* is times right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh then i have to solve it now?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

umm so idk what to do next lol

OpenStudy (anonymous):

\[{8 \over 5} x = 5\] what do you guess to do to solve for X?

OpenStudy (anonymous):



OpenStudy (anonymous):

mukltiply 5?

OpenStudy (anonymous):

Try multiplying both sides by 5 first, what do you get?

OpenStudy (anonymous):

lol cool i was right

OpenStudy (anonymous):

8x+25

OpenStudy (anonymous):

See! You know what to do!

OpenStudy (anonymous):

watch out though, you lost the equals sign and replaced it with a plus sign

OpenStudy (anonymous):

ohh i meant to do = :P then divide by 8? then it comes out as a decimal im sure

OpenStudy (anonymous):

yes it will, but that'll be the length of one side

OpenStudy (anonymous):

oh grr the other side lol

OpenStudy (anonymous):

so what's the length of the other side? what do you think?

OpenStudy (anonymous):

i got 3.125